OpenStudy (anonymous):

Select the factors of x2 − 16x + 64. I know that it is between these two, but I can't figure out which one it is. Please help? (x + 4)(x + 16) (x − 4)(x − 16)

OpenStudy (muzzack):

\[x^2-16x+64 \] Final result : \[ (x - 8)^2\]

OpenStudy (muzzack):

Step by step solution : Step 1 : Simplify x2-16x + 64 Trying to factor by splitting the middle term 1.1 Factoring x2-16x+64 The first term is, x2 its coefficient is 1 . The middle term is, -16x its coefficient is -16 . The last term, "the constant", is +64 Step-1 : Multiply the coefficient of the first term by the constant 1 • 64 = 64 Step-2 : Find two factors of 64 whose sum equals the coefficient of the middle term, which is -16 . -64 + -1 = -65 -32 + -2 = -34 -16 + -4 = -20 -8 + -8 = -16 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -8 and -8 x2 - 8x - 8x - 64 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-8) Add up the last 2 terms, pulling out common factors : 8 • (x-8) Now add up the four terms of step 3 : (x-8) • (x-8) Which is the desired factorization Multiplying Exponential Expressions : 1.2 Multiply (x-8) by (x-8) The rule says : To multiply exponential expressions which have the same base, add up their exponents. In our case, the common base is (x-8) and the exponents are : 1 , as (x-8) is the same number as (x-8)1 and 1 , as (x-8) is the same number as (x-8)1 The product is therefore, (x-8)(1+1) = (x-8)2

OpenStudy (anonymous):

its the second one

OpenStudy (anonymous):

Thank you both for all your help. @Muzzack I really appreciate the step by step, it really cleared things up for me and help me understand how to find the answer. Thank you both again!

OpenStudy (muzzack):

np